Jefferson Lab HALL A Logbook for April 14 - All EXCEPT Automatic Entries

  • Main INDEX

    Wednesday

  • 388892 04/30/14 14:25 R. Michaels mods to ts_scaler*.crl
  • Tuesday

  • 388891 04/29/14 19:58 Fansler/Glor Cold Return
  • 388890 04/29/14 19:36 Fansler/Glor ESR Crash
  • 388889 04/29/14 15:17 fansler Q1 Valve Cal.
  • 388888 04/29/14 12:16 R. Michaels TDC values on Trig. Latch
  • 388887 04/29/14 11:17 R. Michaels short runs to check TS latch
  • 388884 04/29/14 10:53 adaq Trigger-MLU-Settings
  • 388880 04/29/14 10:52 adaq Trigger-MLU-Settings
  • 388877 04/29/14 10:51 adaq Trigger-MLU-Settings
  • 388873 04/29/14 10:50 adaq Trigger-MLU-Settings
  • 388871 04/29/14 10:49 adaq Trigger-MLU-Settings
  • 388868 04/29/14 10:44 adaq Trigger-MLU-Settings
  • 388865 04/29/14 10:35 adaq Trigger-MLU-Settings
  • 388862 04/29/14 10:34 adaq Trigger-MLU-Settings
  • 388859 04/29/14 10:34 adaq Trigger-MLU-Settings
  • 388856 04/29/14 10:33 adaq Trigger-MLU-Settings
  • 388853 04/29/14 10:32 adaq Trigger-MLU-Settings
  • 388850 04/29/14 10:32 adaq Trigger-MLU-Settings
  • 388847 04/29/14 10:30 adaq Trigger-MLU-Settings
  • Monday

  • 388846 04/28/14 15:24 R. Michaels TS latch cabling info
  • 388843 04/28/14 11:28 adaq Trigger-MLU-Settings
  • 388841 04/28/14 00:28 adaq Trigger-MLU-Settings
  • 388839 04/28/14 00:26 adaq Trigger-MLU-Settings
  • Friday

  • 388838 04/25/14 13:53 folts Q1 cooldown
  • Wednesday

  • 388835 04/23/14 15:05 R. Michaels deadtime in datamon from Trig. Super.
  • 388833 04/23/14 14:59 adaq Trigger-MLU-Settings
  • 388831 04/23/14 14:52 adaq Trigger-MLU-Settings
  • 388828 04/23/14 14:51 adaq Trigger-MLU-Settings
  • 388824 04/23/14 14:47 adaq Trigger-MLU-Settings
  • 388822 04/23/14 14:37 adaq Trigger-MLU-Settings
  • 388819 04/23/14 14:37 adaq Trigger-MLU-Settings
  • 388816 04/23/14 14:22 adaq Trigger-MLU-Settings
  • 388813 04/23/14 14:03 adaq Trigger-MLU-Settings
  • 388810 04/23/14 13:57 adaq Trigger-MLU-Settings
  • 388807 04/23/14 13:32 adaq Trigger-MLU-Settings
  • Tuesday

  • 388806 04/22/14 16:21 R. Michaels scaler gating completed
  • 388803 04/22/14 16:16 adaq Trigger-MLU-Settings
  • 388800 04/22/14 16:11 adaq Trigger-MLU-Settings
  • 388798 04/22/14 15:42 adaq Trigger-MLU-Settings
  • 388795 04/22/14 15:22 adaq Trigger-MLU-Settings
  • 388793 04/22/14 15:10 adaq Trigger-MLU-Settings
  • Monday

  • 388790 04/21/14 18:25 adaq Trigger-MLU-Settings
  • 388786 04/21/14 18:23 adaq Trigger-MLU-Settings
  • 388785 04/21/14 14:41 R. Michaels R-HRS scaler gating logic
  • 388782 04/21/14 14:18 adaq Trigger-MLU-Settings
  • 388780 04/21/14 14:01 R. Michaels scalers added to TS readout
  • 388779 04/21/14 13:57 R. Michaels new entries about MLU state
  • 388777 04/21/14 13:53 adaq Trigger-MLU-Settings
  • 388774 04/21/14 13:36 adaq Trigger-MLU-Settings
  • 388771 04/21/14 13:28 adaq Trigger-MLU-Settings
  • 388768 04/21/14 13:26 adaq Settings
  • Friday

    
    

    Thursday

  • 388756 04/17/14 14:40 R. Michaels investigating high-rate T1 on L-HRS
  • 388755 04/17/14 12:39 R. Michaels trigger setup GUI - update
  • Wednesday

    
    

    Tuesday

  • 388734 04/15/14 15:35 R. Michaels MLU: capability to turn off outputs
  • 388712 04/15/14 11:01 R. Michaels renamed hvg directory
  • Friday

    
    

    Thursday

  • 388675 04/10/14 13:43 Kalyan RHRS HV issue fixed
  • 388674 04/10/14 10:35 Kalyan Bad 24 V power supply on RHRS HV crate?
  • 388673 04/10/14 09:43 YANG R-arm HV for FPP and VDC tripped
  • Tuesday

  • 388672 04/08/14 11:54 yxzhao release hallafb3 IP ( hallafb4 is still in use)
  • Monday

  • 388671 04/07/14 11:33 Kalyan Re: Beam energy determination
  • 388670 04/07/14 07:44 Folts trch vesda
  • Friday

  • 388669 04/04/14 17:45 Kalyan Re: Beam energy determination
  • 388668 04/04/14 16:07 Kalyan Beam energy determination
  • 388667 04/04/14 08:01 J. Butler (T Guards Installed
  • Thursday

  • 388666 04/03/14 17:54 A. Camsonne Borrowing 2 IP address for EC SoLID test stand
  • 388665 04/03/14 16:43 Kalyan Thick sieve pattern on left arm
  • 388664 04/03/14 10:21 Higinbotham Spin Calculation
  • 388663 04/03/14 09:18 L. Myers RC summary & End of Run
  • 388662 04/03/14 08:24 georges 3rd April owl shift summary
  • 388661 04/03/14 08:21 georges Target to HOME
  • 388658 04/03/14 08:14 georges online analysis for RHRS run 20372 (bis)
  • 388657 04/03/14 08:13 georges online analysis for RHRS run 20372
  • 388656 04/03/14 08:12 georges online analysis for LHRS run 1407 (bis)
  • 388655 04/03/14 08:11 georges online analysis for LHRS run 1407
  • 388654 04/03/14 07:55 Longwu Online analysis for run 1407
  • 388653 04/03/14 07:38 georges right arm xscaler
  • 388650 04/03/14 07:25 georges xscaler screen (S2m)
  • 388649 04/03/14 07:23 georges xscaler screen
  • 388648 04/03/14 07:23 georges beam is back
  • 388647 04/03/14 07:17 georges beam tripped
  • 388640 04/03/14 07:02 georges target moved to carbon
  • 388639 04/03/14 06:53 georges beam soon
  • 388636 04/03/14 04:36 georges xscaler screen
  • 388635 04/03/14 04:33 georges hall a tools
  • 388634 04/03/14 04:11 georges high voltage screen
  • 388633 04/03/14 04:07 Sasha Moller commissioning result
  • 388632 04/03/14 03:02 georges xscaler screen
  • 388631 04/03/14 01:21 georges xscaler screen
  • 388630 04/03/14 00:12 Mongi Swing shift summary
  • Wednesday

  • 388627 04/02/14 23:56 Mongi xscalers 20368 end RHRS
  • 388626 04/02/14 23:54 Mongi scalers run 1404 end LHRS
  • 388625 04/02/14 21:58 gautam online analysis for run no 20367 RHRS
  • 388624 04/02/14 21:55 gautam Online analysis for run no. 20367 for RHRS
  • 388623 04/02/14 21:38 Mongi Online analysis, run 1403, cosmics continued
  • 388622 04/02/14 21:36 Mongi Online analysis, run 1403, cosmics
  • 388621 04/02/14 21:35 Mongi Online analysis, run 1403, cosmics
  • 388620 04/02/14 21:01 Mongi xscalers RHRS
  • 388619 04/02/14 21:00 Mongi xscalers LHRS
  • 388614 04/02/14 19:32 Sasha Reboot iocsofthabcm1
  • 388613 04/02/14 19:24 Sasha 4.7uA CW beam in the Hall A
  • 388612 04/02/14 19:19 Mongi CW beam in Hall & Moller Group at work
  • 388611 04/02/14 18:07 Mongi Hall A Tools
  • 388610 04/02/14 15:51 desnault Day shift summary April 2nd
  • 388605 04/02/14 14:05 desnault xscaler screens RHRS
  • 388604 04/02/14 14:01 desnault xscaler screens LHRS
  • 388603 04/02/14 13:57 desnault Hall A Tools screen
  • 388602 04/02/14 13:56 desnault HV screen
  • 388601 04/02/14 13:55 N. Hirlinger Main Target Screen
  • 388596 04/02/14 11:18 N. Hirlinger XScalar Window L-HRS
  • 388595 04/02/14 11:17 desnault Hall A Tools screen
  • 388594 04/02/14 11:16 N. Hirlinger XScalar Window R-HRS
  • 388593 04/02/14 11:13 desnault HV Screen
  • 388592 04/02/14 11:09 N. Hirlinger Main Target Screen
  • 388589 04/02/14 09:52 L. Myers RC summary for April 2nd
  • 388584 04/02/14 08:00 N. Israel target screen and charts
  • 388583 04/02/14 08:00 georges owl shift summary (2nd april)
  • 388582 04/02/14 07:58 georges xscaler screen
  • 388581 04/02/14 07:53 georges online analysis for RHRS run 20362
  • 388580 04/02/14 07:53 YANG Gas Cherenkov PMT7 ADC plot for Run 20354
  • 388579 04/02/14 07:52 georges online analysis for RHRS run 20362
  • 388578 04/02/14 07:50 georges online analysis for LHRS run 1399 bis
  • 388577 04/02/14 07:49 georges online analysis for LHRS run 1399
  • 388574 04/02/14 07:01 YANG online analysis Run 20254 (2)
  • 388573 04/02/14 06:58 YANG online analysis Run 20254
  • 388572 04/02/14 06:43 YANG SH, PS plots for Run 20354
  • 388569 04/02/14 06:28 N. Israel target screen and charts
  • 388568 04/02/14 06:28 georges high voltage screen
  • 388567 04/02/14 06:27 george xscaller screen
  • 388566 04/02/14 06:14 N. Israel beam
  • 388559 04/02/14 04:12 N. Israel target screen and charts
  • 388558 04/02/14 03:16 georges xscaller screen
  • 388555 04/02/14 03:12 georges hall A tools screen
  • 388554 04/02/14 02:15 YANG L-arm tdc peak shifted problem still exists
  • 388551 04/02/14 02:06 YANG L-arm FPP HV Scan
  • 388550 04/02/14 01:47 N.Israel target screen and cryocharts
  • 388549 04/02/14 01:40 georges hall A tools screen
  • 388548 04/02/14 01:40 georges high voltage screen
  • 388547 04/02/14 01:38 georges xscaller screen
  • 388544 04/02/14 00:26 Mongi Swing Summary
  • 388541 04/02/14 00:00 Mongi tgt at home
  • Tuesday

  • 388540 04/01/14 23:59 Mongi MCC switching to CW
  • 388536 04/01/14 23:39 Mongi Beam y positions flactuating about nominal
  • 388534 04/01/14 22:45 Mongi xscalers-tune beam, 1uA, RHRS
  • 388533 04/01/14 22:44 Mongi xscalers-tune beam , 1uA, LHRS
  • 388532 04/01/14 22:42 Kalyan 2 inch Pb sieve in for run 1398
  • 388527 04/01/14 22:31 Mongi tgt moved to Carbon tgt
  • 388526 04/01/14 22:30 Mongi Beam back in Hall
  • 388523 04/01/14 21:49 Mongi MCC working on trips,
  • 388522 04/01/14 21:32 Mongi tgt to empty
  • 388521 04/01/14 21:25 Mongi Beam in hall
  • 388520 04/01/14 21:21 Mongi MCC tuning beam into Hall
  • 388517 04/01/14 21:06 Mongi online analysis for RHRS run 20351 continued
  • 388516 04/01/14 21:02 Mongi online analysis for RHRS run 20351
  • 388512 04/01/14 19:28 Mongi tgt moved to home-at MCC request
  • 388511 04/01/14 19:22 Mongi Re: Hall A Tools (R-Dipole stable now)
  • 388508 04/01/14 18:59 Mongi Tgt moved to empty
  • 388507 04/01/14 18:56 Mongi MCC to send beam in approx. 20 minutes
  • 388506 04/01/14 18:48 Mongi HV Screen
  • 388505 04/01/14 18:22 Mongi Hall in Beam Permit
  • 388504 04/01/14 18:00 mongi Hall A Tools
  • 388503 04/01/14 17:59 Mongi xscalers RHRS
  • 388502 04/01/14 17:58 Mongi xscalers LHRS
  • 388501 04/01/14 17:57 Mongi MCC working on vacuum problem, then probably beam afterwards
  • 388499 04/01/14 17:00 Mongi Hall in controlled access
  • 388498 04/01/14 16:51 YANG Announcement in the Hall A difficult to hear clearly
  • 388497 04/01/14 16:48 Mongi Tgt moved to home
  • 388495 04/01/14 16:17 Mongi Sweeping Hall A
  • 388494 04/01/14 15:54 M Jones Restarted gas to GEM chambers
  • 388493 04/01/14 15:46 L. Myers Change: Beam Today
  • 388488 04/01/14 11:19 M Jones Turned off GEm gas bottles while in restricted
  • 388487 04/01/14 11:16 M Jones GEM work yesterday
  • 388486 04/01/14 10:42 Higinbotham Sieve Photo
  • 388485 04/01/14 10:36 YANG For Run1394, the HV of FPP is 1750V
  • 388484 04/01/14 10:33 A. Camsonne BCM plugged back
  • 388483 04/01/14 10:23 A. Camsonne BCM calibration check
  • 388479 04/01/14 09:20 L. Myers 4/1 RC summary
  • 388478 04/01/14 08:29 L. Myers No Beam Today
  • 388477 04/01/14 08:24 R. Michaels more on the plan
  • 388476 04/01/14 08:17 Jack(Techs) Magnet status
  • 388475 04/01/14 08:00 R. Michaels hall status